OpenStudy (anonymous):

For each of the following functions, find the maximum and mimimum values of the function on the circular disk: x^2 + y^2 ≤ 1. Do this by looking at the level curves and gradients. (A) f(x, y) = x + y + 3: maximum value = minimum value = (B) f(x, y) = 3x^2 + 4y^2: maximum value = minimum value = (C) f(x, y) = 3x^2 − 4y^2: maximum value = minimum value =

OpenStudy (anonymous):

(A) ( 1,1) = λ(2x,2y) 2λx = 1 2λy = 1 x = 1/(2λ) y = 1/(2λ) 1/4λ^2 + 1/4λ^2 <= 1 1 + 1 <= 4λ^2 2 <= 4λ^2 4λ^2 - 2 >= 0 2(sqrt(2)λ-1)(sqrt(2)λ+1) = 0 λ = 1/(sqrt(2)), -1/(sqrt(2)) λ = 1/sqrt(2) : x = sqrt(2)/2 y = sqrt(2)/2 λ = -1/sqrt(2) : x = -sqrt(2)/2 y = -sqrt(2)/2 max (sqrt(2)/2,sqrt(2)/2) min (-sqrt(2)/2,-sqrt(2)/2)